<Chapter 18 Homework Exercise 18.46 - Enhanced - with Feedback 20 of 37 Review I Constants I Periodic Table 2Ca(s)02(g)2CaO(s2Ca(s)+O2(g) → 2CaO(s) ΔH∘rxn= -1269.8 kJ; ΔS∘rxn= -364.6 J/K For this problem, assume that all reactants and products are in their standard states. You may want to reference (Pages 811 - 815) Section 18.6 while completing this problem.
Answer #1

given Rx og caes) + 02 (9) 2cao(s) Hex = -1269 kJ osran = - 364-6JIK T = 30°c using this equation sgo - AHO - TOS 14rxn= - 12